PerfectlySoft Corp. is experiencing rapid growth. Dividends are expected to grow at 29 percent per year during the next three years, 13 percent over the following year, and then 4 percent per year thereafter indefinitely. The required return on this stock is 10.04 percent, and the stock currently sells for $62.46 per share. What is the projected dividend (in $) for the coming year? Answer to two decimals, carry intermediate calcs. to four decimals.
Answer #1

P0 = [{D0 * (1 + g1)} / (1 + r)] + [{D0 * (1 + g1)2} / (1 + r)2] + [{D0 * (1 + g1)3} / (1 + r)3] + [{D0 * (1 + g1)3 * (1 + g2)} / (1 + r)4] + [{D0 * (1 + g1)3 * (1 + g2) * (1 + gC)} / {(r - gC) * (1 + r)4}]

$62.46 = [{D0 * (1 + 0.29)} / (1 + 0.1004)] + [{D0 * (1 + 0.29)2} / (1 + 0.1004)2] + [{D0 * (1 + 0.29)3} / (1 + 0.1004)3] + [{D0 * (1 + 0.29)3 * (1 + 0.13)} / (1 + 0.1004)4] + [{D0 * (1 + 0.29)3 * (1 + 0.13) * (1 + 0.04)} / {(0.1004 - 0.04) * (1 + 0.1004)4}]

$62.46 = [1.1723 * D0] + [1.3743 * D0] + [1.6111 * D0] + [1.6544 * D0] + [28.49 * D0]

$62.46 = [34.2988 * D0]

D0 = $62.46 / 34.2988 = $1.82

D1 = D0 * (1 + g1) = $1.82 * (1 + 0.29) = $2.35